OpenStudy (rebeccapink6398):

will medal!!!!!!!! How do you write 78.1 in scientific notation?

OpenStudy (rebeccapink6398):

i want to learn how to do it please :)

OpenStudy (michele_laino):

I write it in this way: \[7.81*10^{1}\] because I shifted the decimal point of one position to left

OpenStudy (rebeccapink6398):

OH! thank you! :D

OpenStudy (michele_laino):

thank you!

OpenStudy (rebeccapink6398):

how about this one? How i do this one? How do you write 5.2 × 10 in standard form?

OpenStudy (rebeccapink6398):



OpenStudy (michele_laino):

no, please \[5.2*10=5.2*10^{1}=52\]